OpenStudy (anonymous):

exponential equation 8^(x^2+(5x/3))=16^(x+(1/2) i meant 8 raise to power x square+ 5x/3=16 raise to power x+ 1/2.

OpenStudy (anonymous):

same gimmick as before write \[8=2^3\] \[16=2^4\] multiply the exponents by 2 and 4 respectively and then set them equal and solve

OpenStudy (anonymous):

*3 and 4 respectively

OpenStudy (anonymous):

\[8^{x^2+\frac{5x}{3}}=16^{x+\frac{1}{2}}\] \[2^{3(x^2+\frac{5x}{3})}=2^{4(x+\frac{1}{2})}\] \[3(x^2+\frac{5x}{3})=4(x+\frac{1}{2})\] I hope you can solve now urself?

OpenStudy (anonymous):

yeah thanks man